In a certain shop , notebooks that normally sell for 59 cents each are on sale at 2 for 99 cents . how much can be saved by purchasing 10 of these notebooks at the sale price ?
A $0.85 B $0.95 C $1.10 D $1.15 E $2.00

Answer by mananth(16946)

normal price = 59 cents
..
sale price = 2 for 99
so 1 costs 99/2 = 49.5 cents
the difference = 59-49.5 = 9.5 cents / book
so for 10 books the difference will be 95 cents
OR $0.95